OpenStudy (anonymous):

MEDAL WILL BE GIVEN OUT !!!! 5. Shannon opens a savings account with $50. Each week after that she deposits $8 into her account. How many weeks does it take her to save $450? (1 point) 56.25 weeks 62.5 weeks 52 weeks 50 weeks

OpenStudy (anonymous):

let w=week 50+8w=450 -50 -50 8w=400 divide both sides by 8 w=50